The cost of 8 adult's tickets and 7 children's tickets is $82.45. The cost of 4 adult's tickets and 9 children's tickets is $65.15. Find the cost of each adult's ticket and each child's ticket. In the answer box below enter the cost of each adult's ticket
Answer by mananth(16946)

x Adult Ticket
y Chuild Ticket
Total tickets
8 x + 7 y = 82.45 .............1

4 x + 9 y = 65.15 .............2
Eliminate y
multiply (1)by 9
Multiply (2) by -7
72 x 63 y = 742.05
-28 x -63 y = -456.05
Add the two equations
44 x = 286.00
/ 44
x = 6.50
plug value of x in (1)
8 x + 7 y = 82
52 + 7 y = 82
7 y = 82 -52
7 y = 30
y = 4.35
Adult Ticket $ 6.50
Child ticket $4.35